### v1

# 00footnotetext:

Abstract
Out-of-Domain (OOD) generalization is the ability of a model trained on one or more domains to generalize to unseen domains. In the ImageNet era of computer vision, evaluation sets for measuring a model’s OOD performance were designed to be strictly OOD with respect to style. However, the emergence of foundation models and expansive web-scale datasets has obfuscated this evaluation process, as datasets cover a broad range of domains and risk test domain contamination. In search of the forgotten domain generalization, we create large-scale datasets subsampled from LAION—LAION-Natural and LAION-Rendition—that are strictly OOD to corresponding ImageNet and DomainNet test sets in terms of style. Training CLIP models on these datasets reveals that a significant portion of their performance is explained by in-domain examples. This indicates that the OOD generalization challenges from the ImageNet era still prevail and that training on web-scale data merely creates the illusion of OOD generalization. Furthermore, through a systematic exploration of combining natural and rendition datasets in varying proportions, we identify optimal mixing ratios for model generalization across these domains. Our datasets and results re-enable meaningful assessment of OOD robustness at scale—a crucial prerequisite for improving model robustness.
Figure 1: Evaluated correctly, CLIP’s OOD performance on renditions drops significantly. A: Models used to be trained on a single domain like natural images from ImageNet (russakovsky2015imagenet) and evaluated for out-of-domain (OOD) generalization on a different domain like renditions from test sets such as ImageNet-R (hendrycks2020many), ImageNet-Sketch (wang2019learning). B: Today, large foundation models like CLIP (radford2021learning) are trained on web-scale datasets such as LAION-400M (schuhmann2021laion) containing images from many domains. Tested on a specific domain like renditions, CLIP exhibits unprecedented performance and appears robust. C: We subsample from a deduplicated LAION-400M (abbas2023semdedup) to obtain LAION-Natural, a web-scale dataset containing only natural images, which re-enables a meaningful assessment of CLIP’s generalization performance to renditions. D: CLIP trained on LAION-Natural performs noticeably poorer on renditions, suggesting that its OOD performance has been previously overestimated. The models are evaluated on refined test datasets containing samples only from their intended domains.
1 Introduction
Foundation models have revolutionized our world, demonstrating remarkable capabilities in solving grade school math problems, writing creative essays, generating stunning images, and comprehending visual content openai2023gpt4; ChatGPT; ramesh2022hierarchical. One notable example is CLIP (radford2021learning), a vision-language model pretrained on a vast dataset of image-text pairs, which forms the backbone of numerous other foundation models (ramesh2022hierarchical; liu2023visualinstructiontuning). CLIP has achieved unprecedented performance across a wide range of benchmarks spanning many domains—a sharp contrast to models from the ImageNet era, which struggled to generalize from a training domain mostly consisting of natural photographs to stylistically different domains such as ImageNet-Sketch (wang2019learning), ImageNet-R (hendrycks2020many), and DomainNet (peng2019moment).
Domains, while not always rigorously defined, emerge from collecting data from specific sources and conditions. Some domains, like natural images or renditions, are better delineated, allowing the creation of datasets like the ones mentioned above. Out-of-domain (OOD) generalization refers to a model’s ability to perform well on data from domains other than its training domain(s) (wang2022generalizingunseendomainssurvey). In this work, we collectively refer to the domain represented by ImageNet-Sketch, ImageNet-R, DomainNet-Painting, DomainNet-Clipart, DomainNet-Sketch, and DomainNet-Quickdraw as the rendition domain, since it contains images that are renditions of natural objects and scenes. Generalization to the rendition domain (especially OOD) is crucial for aligning models with human perception, as humans can interpret abstract visual renditions, while machines tend to rely heavily on textural cues (hendrycks2020many; geirhos2018imagenet).
CLIP’s strong performance in several domains, including renditions, is attributed to its vast training distribution, rather than its contrastive learning objective, language supervision, or dataset size (fang2022data). However, fang2022data do not specify what characteristics of the training distribution drive this performance. CLIP could be learning more robust representations due to the diversity of natural images in its training set—or it may simply have been exposed to many datapoints from the (assumed to be OOD) test domains during training. Indeed, mayilvahanan2024 revealed that CLIP’s training data contains exact or near duplicates of samples of many OOD datasets. Yet, they showed that CLIP still generalizes well when this sample contamination is corrected. However, their analysis failed to account for domain contamination.
In contrast to sample contamination, domain contamination does not care about duplicates of specific datapoints but instead checks whether crucial aspects of a test domain are included in the training domain, e.g., by including images with different content but similar style to test samples. For example, after the correction by mayilvahanan2024, although not duplicates, many other rendition images remained in CLIP’s training set (refer to Tab. LABEL:tab:domain_composition).s from the test domain. This leads us to our central question:
To what extent does domain contamination explain CLIP’s performance on renditions?
We address the central question with the following contributions:
[…]
  • Refining the Evaluation of CLIP’s OOD Performance: Using LAION-Natural, we demonstrate that CLIP trained only on natural images significantly underperforms on rendition domain shifts (LABEL:sec:laion_nat_v_random). This suggests that its original success stems from domain contamination, not from an intrinsic OOD generalization ability (see Fig. 1 for a summary).
  • Investigating Domain Mixing and Scaling Effects: Our single-domain datasets enable analyzing the effects of training on controlled mixtures of natural and rendition images across scales (LABEL:sec:laion_mix). We identify the optimal mixing ratio for the best overall performance and show the degree to which training on one domain enables some generalization to the other.
Through this work, we aim to shed light on the limitations of foundation models like CLIP in handling OOD generalization and provide valuable datasets and tools to the community for further exploration. Fig. 1 illustrates our core methodology.
2 Related Work
Measuring the OOD Generalization of CLIP Models
We aim to understand the OOD generalization capabilities of CLIP from a data-centric viewpoint. While multi-modal training with rich language captions does seem to contribute to robustness against distribution shifts (xue2024understanding), fang2022data demonstrated that the nature of CLIP’s training distribution (as opposed to its mere size, its specific training objective, or natural language supervision) causes strong performance on various distribution shifts. However, it is unclear what aspects of the data distribution drive the robustness gains. mayilvahanan2024 remove images highly similar to the test sets to show that data contamination and high perceptual similarity between training and test data do not explain generalization performance. While their data pruning technique removes some samples from LAION-400M that lie outside the natural image domain, they do not address domain generalization: They only account for the part of a domain covered by existing test sets and give no guarantee that all images of a given domain were removed. In another line of work, nguyen2022quality discover that a model’s effective robustness (fang2022data; taori2020measuring) on a test set interpolates when training data is compiled from various sources. However, they only consider mixing datasets that each cover multiple domains. In this work, we take their analysis further and show how mixing two data sources from distinct domains interpolates the effective robustness on those domains. Our study’s title is inspired by gulrajani2020search, who studied generalization from multiple distinct source domains. In contrast, we focus on generalization from single or mixed source domains to unseen domains.
Domain Classification
The primary goal of our work necessitates creating web-scale datasets of different domains. This entails building a robust domain classifier that can reliably distinguish natural images from renditions. This task can be regarded as classifying the style of an image, which Gatys2015ANA proposed to measure using Gram matrices and which has been widely explored since then (sandoval2019; menis-mastromichalakis2020; sandoval2018; joshi2020; garcia2018; chu2018; bai2021). More recently, cohen2024ask use a fine-tuned CLIP model from OpenCLIP (ilharco_gabriel_2021_5143773) to distinguish between ImageNet and test sets with a domain shift, such as ImageNet-Sketch, ImageNet-R, and ImageNet-V2 (recht2019imagenet). wang2023evaluating and somepalli2024measuring develop a dataset classifier using a backbone trained by self-supervised learning and classification through retrieval via a database. liu2024decade report high performance when training image classifiers to distinguish between different large-scale and diverse datasets.
3 Constructing Clean Single-Domain Datasets
To answer our central question—how much of CLIP’s performance on renditions can be explained by domain contamination—we must filter out datapoints from specific domains within web-scale datasets. Similar to how ImageNet is compared to ImageNet-Sketch and ImageNet-R, or how DomainNet-Real is compared to DomainNet-Sketch (Quickdraw, Infograph, Clipart, and Painting), we aim to create clean natural and rendition datasets from LAION by building a domain classifier to distinguish between these domains.
To build a robust domain classifier, we first create a labeled dataset where each class represents a distinct domain. The labeling process is outlined in Sec. 3.1, and we explore different ways to build a domain classifier in Sec. 3.2. Further, in LABEL:sec:testing_domain_classifier, we employ the best-performing classifiers to analyze the composition of different training and test sets and finally use it to subsample LAION-Natural and LAION-Rendition in LABEL:sec:subsampling_datasets.
LAION-200M
For the remainder of this work, we substitute LAION-400M with LAION-200M, which we obtain by de-duplicating LAION-400M based on perceptual similarity as introduced by abbas2023semdedup. Both abbas2023semdedup and mayilvahanan2024 demonstrate that CLIP trained on LAION-200M obtains comparable downstream performance while greatly reducing the computational burden of training models from scratch and analyzing the dataset.
3.1 Labeling
LAION-200M contains diverse images from a multitude of sources. The images vary from naturally occurring to synthetically generated. We encourage the reader to glance at LABEL:fig:laion_random to get a sense of the dataset and the difficulty of determining the domain of each image. We aim to classify these images mainly as natural or renditions. We also add an extra ambiguous class for images with elements of both domains, images with elements of neither, and edge cases.
We manually label images based on a labeling handbook derived from analyzing the existing OOD test sets, which we outline in LABEL:sec:appx_labeling. In general, we adopt a texture-centric approach to distinguish renditions of a scene or object from their natural depictions. That is, depictions where fine-grained texture information is preserved are generally considered natural, while depictions with simplified or flat textures are considered renditions. Fig. 2 illustrates this demarcation on samples from LAION-200M, ImageNet test sets, and DomainNet test sets.
Figure 2: Labeled natural, ambiguous, and rendition samples from different datasets. Natural images are photos or high-quality renders with minor filters that preserve fine-grained textures, while renditions are typically sketches, paintings, or graphics with flat or simplified textures. Images with elements of both, such as collages or natural images with large stylized elements, and images mainly containing text are labeled as ambiguous.
To further ease the labeling procedure, we first build a rough binary classifier by fine-tuning CLIP ViT-L/14 with a linear readout to differentiate between some of the natural ImageNet and DomainNet test sets (namely, ImageNet-Val, ObjectNet (barbu2019objectnet), ImageNet-V2, ImageNet-A (hendrycks2021natural), and DomainNet-Real) and rendition test sets (namely, ImageNet-Sketch, ImageNet-R, DomainNet-Painting, DomainNet-Sketch, and DomainNet-Clipart). We use this classifier to roughly pre-label samples before they are annotated by a human. The annotator verifies and potentially updates the labels for $25$ images at a time (see LABEL:fig:labeling_setup).
Overall, we label 19 0001900019\,00019 000 random images from LAION-200M and $1000$ images from each of the ImageNet and DomainNet test sets (12 0001200012\,00012 000 in total). Notably, almost all ImageNet and DomainNet test sets usually assumed to contain only images of a single domain exhibit some domain contamination. We discuss this in detail in LABEL:sec:testing_domain_classifier. LABEL:tab:labeling_nos contains a detailed breakdown of labels for each dataset. We show more samples grouped by domain for each dataset in LABEL:fig:imagenet_a and LABEL:fig:dn_sketch_1k_visualization.
3.2 Training and Choosing the Domain Classifier
With the domain-labeled dataset, we can train a domain classifier to partition all of LAION-200M into natural images, renditions, or ambiguous images. Since we aim to obtain datasets containing only images from a single domain, we need a domain classifier that is as precise as possible. To this end, we train classifiers on 13 0001300013\,00013 000 labeled LAION-200M images, retaining $3000$ samples each for a validation and test set. From the domain classification literature discussed in Sec. 2, we evaluate four methods with publicly available code. All methods build on CLIP ViT-L/14 pretrained on LAION-2B, which we choose for its balance between accuracy and inference speed. For brevity, we present the two methods we finally employ here, and refer the reader to  LABEL:sec:appx_dc_full for a detailed description, results, and comparisons with other approaches.
Density Ratios
cohenwang2024ask aim to estimate the probability that a given sample is drawn from a reference distribution $p_{\text{ref}}$. Since high dimensional density estimation is challenging, they build a classifier to distinguish between a reference and a shifted distribution and compute the density ratio $\frac{p_{\text{ref}}}{p_{\text{shifted}}}$ which they threshold at $0.2$ to classify a given sample. We deploy their method unchanged to our task. We obtain two binary classifiers, DR-N and DR-R, that distinguish natural from non-natural samples and renditions from non-renditions, respectively.
Fine-Tuning
We fine-tune pretrained CLIP ViT-L/14’s image encoder with a randomly initialized linear readout on the training dataset to obtain a ternary classifier, dubbed FT.
[…]
Table 1: We chose the best natural classfier and the best rendition classifier between a binary classifier based on (DR) (cohenwang2024ask) and a ternary classifier using a linear readout based on fine-tuned CLIP model (FT). All models use CLIP ViT-L/14 pretrained on LAION-2B. We report precision and recall for the natural class (top) and rendition class (bottom) on ImageNet (IN) and DomainNet (DN) test sets and average performance across all test sets. For each class, we select the classifier with the highest validation-recall.
[…]

### v2

@@
 1 Introduction
-Foundation models have revolutionized our world, demonstrating remarkable capabilities in solving grade school math problems, writing creative essays, generating stunning images, and comprehending visual content openai2023gpt4; ChatGPT; ramesh2022hierarchical.
+Foundation models have revolutionized our world, demonstrating remarkable capabilities in solving grade school math problems, writing creative essays, generating stunning images, and comprehending visual content (openai2023gpt4; ChatGPT; ramesh2022hierarchical).
 One notable example is CLIP (radford2021learning), a vision-language model pretrained on a vast dataset of image-text pairs, which forms the backbone of numerous other foundation models (ramesh2022hierarchical; liu2023visualinstructiontuning).
 CLIP has achieved unprecedented performance across a wide range of benchmarks spanning many domains—a sharp contrast to models from the ImageNet era, which struggled to generalize from a training domain mostly consisting of natural photographs to stylistically different domains such as ImageNet-Sketch (wang2019learning), ImageNet-R (hendrycks2020many), and DomainNet (peng2019moment).
-Domains, while not always rigorously defined, emerge from collecting data from specific sources and conditions.
+Domains, while often challenging to quantify in practice (bendavid), emerge from collecting data from specific sources and conditions.
 Some domains, like natural images or renditions, are better delineated, allowing the creation of datasets like the ones mentioned above.
@@
 However, their analysis failed to account for domain contamination.
-In contrast to sample contamination, domain contamination does not care about duplicates of specific datapoints but instead checks whether crucial aspects of a test domain are included in the training domain, e.g., by including images with different content but similar style to test samples.
-For example, after the correction by mayilvahanan2024, although not duplicates, many other rendition images remained in CLIP’s training set (refer to Tab.
-LABEL:tab:domain_composition).s from the test domain.
+In contrast to sample contamination, domain contamination does not focus on duplicates of specific datapoints but rather examines whether critical aspects of a test domain are present in the training domain, such as images with different content but similar style to test samples.
+For example, after the correction by mayilvahanan2024, many other rendition images, while not duplicates, remained in CLIP’s training set (refer to Tab.
+LABEL:tab:domain_composition).
+Prior works often assume that CLIP is capable of generalizing OOD (radford2021learning; abbasi2024decipheringrolerepresentationdisentanglement; nguyen2024saftoutofdistributiongeneralizationfinetuning; fang2022data; li2023distillinglargevisionlanguagemodel; shu2023clipoodgeneralizingclipoutofdistributions); however, it remains unclear whether this is truly the case or if its performance is primarily driven by training on images from the test domain.
 This leads us to our central question:
@@
 In contrast, we focus on generalization from single or mixed source domains to unseen domains.
+Overall, we aim for our work to be a valuable addition to the literature on OOD generalization (liu2023outofdistributiongeneralizationsurvey; koh2021wildsbenchmarkinthewilddistribution; madan2021cnnsgeneralizeoutofdistributioncategoryviewpoint; gulrajani2020search; madan2022what; arjovsky_invariant_2020; arjovsky_out_2021).
 Domain Classification